1. What is the pH of a 0.34 M solution of (weak acid) HF if the Ka = 6.8x10-4

?


2. If the pH of a weak acid solution is 9.5 and the original concentration of base was 0.30 M

A. what is the pOH?

B. what is the concentration of OH-

?


C. what is the equilibrium concentration of the base?


D. what is the Kb of the base?


3. Find the pH of a 0.325 M (weak acid) acetic acid solution. Ka = 1.8 x 10-5

.


4. Find the pH of a 0.056 M (weak acid)propionic acid solution (Ka = 1.4 x 10-5

).


5. Find the pH of a 0.075 M solution of (weak acid) formic acid. The acid dissociation

constant (Ka) for formic acid is 1.8 x 10-4

.


6. Find the pH of a 0.15 M solution of (weak base) ammonia, NH3. Kb = 1.8 X 10-5


7. Find the pH of a 0.600 M solution of (weak base) methylamine CH3NH2. Kb = 4.4 x 10–4

.


8. If the pH of (weak acid) HC3H5O2 is 4.2 and the Ka = 1.34x10-5

A. what is the equilibrium concentration of HC3H5O2?


B. what was the initial concentration of HC3H5O before dissociation?


1
Expert's answer
2021-05-31T02:26:07-0400

1) pH=0.34M*6.8^10-4=2.312^10-4


2) A. This is a measure of the acidity of the ionic OH-

B. pH>7 (since this is the foundation)

С. The equilibrium concentration is equal to the product of the concentration and the amount of the base substance


3)pH=0.325M*1.8^10-5=0.585


4)pH=0.056M*1.4^10-5=0.0784


5)pH=0.075M*1.8^10-4=0.0135


6)pH=0.15M*1.8^10-5=0.27


7)pH=0.6M*4.4^10-4=2.64


8)C=4.2/1.34^10-5=3.134